the current price is $1.00 per unit and quantity demand is 2500 per unit based on empirical studies you know that price elasticity of demand for product is 0.5 if the price is increase to $2.00 per unit what is predicted % in quantity demand and total expenditure increase or decrease
Answer #1

Price elasticity of demand = 0.5

Percentage increase in price = [(2-1)/(1+2)/2]100 =66.67%

Price elasticity of demand= (% change in quantity demanded/% change in price)

0.5 = (% decrease in quantity demanded)/66.67

0.5(66.67)= % decrease in quantity demanded

So,the predicted % decrease in quantity demanded = 33.33%

Because elasticity of demand is inelastic , so increase in price would increase total expenditure.

Here, when P=$1 , TR= (1)(2500)= $2500

New quantity =2500- (2500)(33.33%) =1666.75

And When P=$2 , TR=(2)(1666.75)= $3333.5